# N-Mapping: Normal Guided Neural Non-Projective Signed Distance Fields for Large-scale 3D Mapping

### v1

@@ Abstract
Accurate and dense mapping in large-scale environments is essential for various robot applications. Recently, implicit neural signed distance fields (SDFs) have shown promising advances in this task. However, most existing approaches employ projective distances from range data as SDF supervision, introducing approximation errors and thus degrading the mapping quality. To address this problem, we introduce N-Mapping, an implicit neural mapping system featuring normal-guided neural non-projective signed distance fields. Specifically, we directly sample points along the surface normal, instead of the ray, to obtain more accurate non-projective distance values from range data. Then these distance values are used as supervision to train the implicit map. For large-scale mapping, we apply a voxel-oriented sliding window mechanism to alleviate the forgetting issue with a bounded memory footprint. Besides, considering the uneven distribution of measured point clouds, a hierarchical sampling strategy is designed to improve training efficiency. Experiments demonstrate that our method effectively mitigates SDF approximation errors and achieves state-of-the-art mapping quality compared to existing approaches.
I INTRODUCTION
@@
Building accurate and dense maps within large-scale environments is crucial for various robot applications, such as autonomous driving. The signed distance field (SDF) has been an important map representation employed to accomplish this task, due to its advantages in characterizing geometric information and compatibility with downstream tasks [1]. Unfortunately, traditional SDF mapping methods [2, 3, 4] have faced longstanding challenges in striking a balance between computational consumption and mapping accuracy. The primary reasons that limit mapping accuracy are twofold. The first is the discretization error because the space is divided into voxel grids that store SDF values, and this error grows as the grid size increases. The second is the approximation error introduced by using projective distance, i.e. the distance along range sensor rays to the measured surface, to estimate SDF values. This can easily lead to an overestimation of the actual distance to the nearest surface.
@@
Recently, with the introduction of DeepSDF [5], NeRF [6] and Occupancy Networks [7], neural implicit representations have shown promising potential in modeling 3D scenes. Neural network-based distance fields exhibit continuity, enabling them to overcome resolution limitations and achieve high-fidelity reconstructions. However, learning accurate distance fields from raw range sensor data is not trivial, as it is challenging to obtain true distance supervision. Most current methods [8, 9, 10, 11, 12] still employ projective distance to approximate the ground-truth SDF supervision for training efficiency, while neglecting the associated systematic errors. Several works [13, 14] leverage surface normals to correct the projective distance and reduce the approximation error. Nonetheless, the former [13] focuses solely on large planar surfaces, limiting its applicability. The latter [14], on the other hand, lacks training stability as it approximates the normal direction using the unstable gradients of neural networks during training. Currently, it is vital to develop a practical non-projective SDF mapping system to overcome these limitations.
@@
To this end, we propose N-Mapping, a large-scale mapping approach with normal guided neural non-projective signed distance fields. We observe that, for any point around the surface, the normal generally provides the direction to the nearest surface point. Therefore, our method directly samples points and corresponding distance values along the normal direction near the surface. Such sampled SDF labels tend to be close to the ground truth, leading to improved mapping quality. For the implicit map, our method utilizes octree-based hierarchical sparse voxels to store optimizable feature vectors and a shallow MLP to decode queried local features into SDF values.
@@ II RELATED WORK
Various scene representations have been explored for 3D dense mapping, including occupancy grids [15, 16], meshes [17, 18, 19] and surfels [20, 21]. In addition to these, SDFs have gained significant popularity in many robotic tasks such as planning [22, 2], localization [23, 14], and 3D mapping [4, 3]. This is because SDFs inherently store the signed distance value to the nearest surface and can achieve sub-voxel resolutions in reconstruction. As a milestone of using truncated signed distance function (TSDF) as map representation, KinectFusion [4] uses depth images as input and introduces the volumetric integration algorithm to enable real-time mapping. Subsequent studies follow the manner of TSDF integration and work on enhancing efficiency [2, 3], accuracy [24], and the capability to handle dynamic [25] and large-scale environments [26]. However, constrained by their explicit representations, these methods require substantial memory resources to achieve high resolutions.
@@
Recently, research on implicit neural representations has made significant progress. Works like DeepSDF [5], Occupancy Networks [7], and NeRF [6] use implicit neural networks to represent 3D objects and scenes, overcoming the limitations of traditional explicit representations. Inspired by them, [27, 28, 8] achieve incremental neural mapping by continuously training an MLP to represent the environment with sequentially input data. Considering the limited capacity of a single MLP, NICE-SLAM [29] combines dense grids to store optimizable local features and employs shallow MLPs to decode hidden information. This combination allows for more accurate and faster reconstruction but dense grids are not memory-efficient. Subsequent works tackle this issue by employing various data structures such as sparse octree [11, 10], hash encoding [30], and neural points [31]. In our approach, we adopt a sparse octree to store multi-layer local features, as done in SHINE-Mapping [11].
@@
For training SDF with range data, recent studies commonly use projective distance [8, 9, 10, 11, 12], i.e. the signed distance from a sampled point along the ray to the endpoint, as the supervision signal. This SDF approximation can lead to faster convergence while it also introduces errors, especially when the incidence angle of the ray is small. To obtain more accurate non-projective SDF, Voxfield [24] performs normals integration for each voxel and uses the fused normal to correct the distance value. Similarly, NeRF-LOAM [13] applies projective distance correction with normals specifically for ground points, and LocUDF [14] uses the gradient of the distance field to rectify SDF labels. Nevertheless, NeRF-LOAM only considers the ground plane, and LocUDF faces training instability due to circular dependency. Additionally, [28, 32] compute the distance between sampled points and the nearest measured surface points as the SDF label. However, such methods are sensitive to noise and rely on pre-built dense point clouds [32]. In our method, accurate SDF labels can be easily obtained by directly sampling points along the normal direction.
@@
III N-Mapping
In this paper, we present N-Mapping, a framework designed for large-scale and high-quality 3D mapping. as outlined in Figure 1. Given sequentially input points and normals from LiDAR or RGB-D cameras with associated poses, we obtain accurate non-projective SDF labels, along with corresponding sampled points, through normal-guided sampling (Section III-B). These training pairs are then maintained in our voxel-oriented sliding window (Section III-C1). For subsequent training, by employing a hierarchical sampling strategy (Section III-C2), we select a batch of training pairs at each iteration to optimize our octree-based implicit map (Section III-A2). Finally, we visualize and evaluate the mapping results as 3D meshes generated through the marching cubes algorithm [17].
@@ III-A1 Signed Distance Fields
A signed distance field represents the 3D scene by assigning a signed distance to each point in space relative to the closest surface. The sign of the distance indicates whether a point is inside or outside the surface. The signed distance function can be defined as: , which maps a 3D coordinate to the signed distance value . It is worth noting that on the surface, the gradient of w.r.t. is equal to the surface normal: (or depending on the normal’s orientation). This implies that normal priors can provide informative guidance for building an SDF map.
@@ III-B1 Normal Guided Non-projective Distance
We observe that along the negative normal direction, the sampled points direct toward the closest surface, aligning with the gradient direction of the distance field. Therefore, we propose to directly sample points along the normal direction and use the signed distance from the sampled point to the surface point as the supervision:
@@ IV-A3 Implementation Details
Our method takes as input point-wise normals which are estimated using Open3D [36] with k-nearest neighbors set to 20. For implicit map representation, feature vectors are stored in the lowest three levels of the octree, each with a length of 8. The leaf voxel size is set to . Our shared MLP decoder contains 2 fully-connected layers with ReLU activations, and each layer has 32 hidden units. We sample 3 points near the surface and another 3 in free space. The noise hyperparameter is 0.1 and the truncation distance is . For training, we set and . All experiments are conducted on a desktop PC with a 3.7 GHz Intel i9-10900X CPU and an NVIDIA RTX3090 GPU.
@@ IV-D Scalable Incremental Mapping
IV-E Runtime Analysis
We choose Maicity-01 and KITTI-07 to evaluate the computational efficiency of our method. Both our method and SHINE-Mapping perform 50 iterations per frame, and the voxel size is set to 0.2m. As shown in Table IV, although our method takes longer for training pairs allocation in the preprocessing stage, it avoids the time-consuming regularization update during training, resulting in better efficiency. Nevertheless, our method still cannot run in real time. It can be further sped up by the optimized C++ implementation, faster map management [39] and more efficient implicit representations [12, 30].
@@ V CONCLUSION
We introduce N-Mapping, a novel implicit neural mapping approach designed for large-scale environments. Our method does not rely on projective distances along the ray but opts to generate more accurate SDF values along normal directions. These non-projective SDF values serve as supervision signals for training the implicit map, resulting in superior reconstruction performance. Additionally, we implement a voxel-oriented training strategy, incorporating a sliding window for bounded memory consumption and a hierarchical sampling method to enhance training efficiency. Experiments show that our mapping results are more accurate and complete than existing methods, demonstrating the effectiveness of our designs.

### v2

# N3-Mapping: Normal Guided Neural Non-Projective Signed Distance Fields for Large-scale 3D Mapping

@@ Abstract
Accurate and dense mapping in large-scale environments is essential for various robot applications. Recently, implicit neural signed distance fields (SDFs) have shown promising advances in this task. However, most existing approaches employ projective distances from range data as SDF supervision, introducing approximation errors and thus degrading the mapping quality. To address this problem, we introduce N3-Mapping, an implicit neural mapping system featuring normal-guided neural non-projective signed distance fields. Specifically, we directly sample points along the surface normal, instead of the ray, to obtain more accurate non-projective distance values from range data. Then these distance values are used as supervision to train the implicit map. For large-scale mapping, we apply a voxel-oriented sliding window mechanism to alleviate the forgetting issue with a bounded memory footprint. Besides, considering the uneven distribution of measured point clouds, a hierarchical sampling strategy is designed to improve training efficiency. Experiments demonstrate that our method effectively mitigates SDF approximation errors and achieves state-of-the-art mapping quality compared to existing approaches. The code will be released at https://github.com/tiev-tongji/N3-Mapping.
Index Terms:
Mapping; SLAM; Implicit Neural RepresentationsI INTRODUCTION
@@
Building accurate and dense maps within large-scale environments is crucial for various robot applications, such as autonomous driving. The signed distance field (SDF) has been an important map representation employed to accomplish this task, due to its advantages in characterizing geometric information and compatibility with downstream tasks [1]. Unfortunately, traditional SDF mapping methods [2, 3, 4] have faced two longstanding challenges. The first is the trade-off between mapping accuracy and memory consumption. The second is the approximation error introduced by using projective distance, i.e. the distance along range sensor rays to the measured surface, to estimate SDF values. This can easily lead to an overestimation of the actual distance to the nearest surface.
@@
Recently, neural implicit representations [5, 6, 7] have shown promising potential in modeling 3D scenes. Neural network-based distance fields exhibit continuity, enabling them to overcome resolution limitations and achieve high-fidelity reconstructions. However, learning accurate distance fields from raw range sensor data is not trivial, as it is challenging to obtain true distance supervision. Most current methods [8, 9, 10, 11, 12] still employ projective distance to approximate the ground-truth SDF supervision for training efficiency, while neglecting the associated systematic errors. Several works [13, 14] leverage surface normals to correct the projective distance and reduce the approximation error. Nonetheless, the former [13] focuses solely on large planar surfaces, limiting its applicability. The latter [14], on the other hand, lacks training stability as it approximates the normal direction using the unstable gradients of neural networks during training. Currently, it is vital to develop a practical non-projective SDF mapping system to overcome these limitations.
@@
To this end, we propose N3-Mapping, a large-scale mapping approach with normal guided neural non-projective signed distance fields. We observe that, for any point around the surface, the normal generally provides the direction to the nearest surface point. Therefore, our method directly samples points and corresponding distance values along the normal direction near the surface. Such sampled SDF labels tend to be close to the ground truth, leading to improved mapping quality. For the implicit map, our method utilizes octree-based hierarchical sparse voxels to store optimizable feature vectors and a shallow MLP to decode queried local features into SDF values.
@@ II RELATED WORK
Various scene representations have been explored for 3D dense mapping, including occupancy grids [15], meshes [16] and surfels [17]. In addition to these, SDFs have gained significant popularity in many robotic tasks such as planning [18, 2], localization [19, 14], and 3D mapping [4, 3]. As a milestone of using truncated signed distance function (TSDF) as map representation, KinectFusion [4] uses depth images as input and introduces the volumetric integration algorithm to enable real-time mapping. Subsequent studies follow the manner of TSDF integration and work on enhancing efficiency [2, 3], accuracy [20], and the capability to handle dynamic [21] and large-scale environments [22]. However, constrained by their explicit representations, these methods require substantial memory resources to achieve high resolutions.
@@
Recently, research on implicit neural representations has made significant progress. Seminal works [5, 6, 7] use implicit neural networks to represent 3D objects and scenes, overcoming the limitations of traditional explicit representations. Inspired by them, [23, 24, 8] achieve incremental neural mapping by continuously training an MLP to represent the environment with sequentially input data. Considering the limited capacity of a single MLP, NICE-SLAM [25] combines dense grids to store optimizable local features and employs shallow MLPs to decode hidden information. This combination allows for more accurate and faster reconstruction but dense grids are not memory-efficient. Subsequent works tackle this issue by employing various data structures such as sparse octree [11, 10], hash encoding [26], and neural points [27]. In our approach, we adopt a sparse octree to store multi-layer local features, as done in SHINE-Mapping [11].
@@
For training SDF with range data, recent studies [8, 9, 10, 11, 12] commonly use projective distance, i.e. the signed distance from a sampled point along the ray to the endpoint, as the supervision signal. This SDF approximation can lead to faster convergence while it also introduces errors, especially when the incidence angle of the ray is small. To obtain more accurate non-projective SDF, Voxfield [20] performs normals integration for each voxel and uses the fused normal to correct the distance value. Similarly, NeRF-LOAM [13] applies projective distance correction with normals specifically for ground points, and LocNDF [14] uses the gradient of the distance field to rectify SDF labels. Nevertheless, NeRF-LOAM only considers the ground plane, and LocNDF faces training instability due to circular dependency. Additionally, [24, 28] compute the distance between sampled points and the nearest measured surface points as the SDF label. However, such methods are sensitive to noise and rely on pre-built dense point clouds [28]. In our method, accurate SDF labels can be easily obtained by directly sampling points along the normal direction.
@@
III N3-Mapping
In this paper, we present N3-Mapping, a framework designed for large-scale and high-quality 3D mapping. as outlined in Figure 1. Given sequentially input points and normals from LiDAR or RGB-D cameras with associated poses, we obtain accurate non-projective SDF labels, along with corresponding sampled points, through normal-guided sampling (Section III-B). These training pairs are then maintained in our voxel-oriented sliding window (Section III-C1). For subsequent training, by employing a hierarchical sampling strategy (Section III-C2), we select a batch of training pairs at each iteration to optimize our octree-based implicit map (Section III-A2). Finally, we visualize and evaluate the mapping results as 3D meshes generated through the marching cubes algorithm [16].
@@ III-A1 Signed Distance Fields
A signed distance field represents the 3D scene by assigning a signed distance to each point in space relative to the closest surface. The sign of the distance indicates whether a point is inside or outside the surface. The signed distance function can be defined as: , which maps a 3D coordinate to the signed distance value .
Notably, the gradient of SDF points towards the closest surface and on the surface the gradient is equal to the surface normal: (or depending on the normal’s orientation). This implies that normal priors can provide informative guidance for neural SDF learning.
@@ III-B1 Normal Guided Non-projective Distance
As discussed in Section III-A1, along the negative normal direction, the sampled points direct toward the closest surface, aligning with the gradient direction of the distance field. Therefore, we propose to directly sample points along the normal direction and use the signed distance from the sampled point to the surface point as the supervision:
@@ IV-A3 Implementation Details
Our method takes as input point-wise normals which are estimated using Open3D [32] with k-nearest neighbors set to 20. For implicit map representation, feature vectors are stored in the lowest three levels of the octree, each with a length of 8. The leaf voxel size is set to and the truncation distance is . We sample 3 points near the surface and another 3 in free space. Our shared MLP decoder contains 2 fully-connected layers with ReLU activations, and each layer has 32 hidden units. For training, we set , and the sharpness parameter . All experiments are conducted on a desktop PC with a 3.7 GHz Intel i9-10900X CPU and an NVIDIA RTX3090 GPU.
@@ IV-D Scalable Incremental Mapping
IV-E Robustness Analysis
We test our method on the Neural RGBD dataset [8] to evaluate its robustness in different environments. Here we set the leaf voxel size to 4 and truncation distance to 3 to accommodate indoor environments. Figure 8 demonstrates that our approach effectively captures intricate details of thin and complex structures, such as the chair and the crate. As a reference, the results obtained by using projective distance exhibit a “bulging effect” and suffer from noticeable artifacts.
We also evaluate the impact of truncation distance on reconstruction quality. As indicated in Figure 9, our method achieves lower C-L1 than those using projective SDF. However, our method also exhibits a slightly higher sensitivity to the truncation distance in the Staircase scenario with complex geometries. This is because, in such cases, normal guided sampled points can easily interfere with other nearby surfaces, resulting in poorer reconstruction. When applied in structurally simpler urban scenes (Maicity), our method demonstrates increased robustness to the choice of the truncation interval. It is also worth noting that an excessively small truncation distance can compromise the completeness of the reconstruction, leading to a significant increase in overall error.
Lastly, we take a close look at the impact of moving obstacles on the mapping results, using the Newer College dataset as an example. As depicted in Figure 10, during the mapping process, moving objects like the pedestrians leave artifacts along their motion trajectories. However, thanks to our free space sampling design, these artifacts can be gradually suppressed and eliminated, while static objects such as the van are retained. Consequently, the final mapping results are complete and clean.
IV-F Runtime Analysis
We choose Maicity-01 and KITTI-07 to compare the runtime in Table IV. Although our method takes longer for training pairs allocation in the preprocessing stage, it avoids the time-consuming regularization update during training, resulting in better efficiency. Currently, our method is unable to meet real-time requirements and performs incremental mapping offline. It can be further sped up by the optimized C++ implementation and more efficient implicit representations [12, 26].
@@ V CONCLUSION
We introduce N3-Mapping, a novel large-scale mapping approach using normal guided neural non-projective SDFs. Experiments demonstrate that our method overcomes the limitations associated with projective distance and achieves more accurate and complete reconstruction. Additionally, our voxel-oriented training strategy not only enables efficient incremental mapping without the forgetting issue but also ensures bounded memory consumption. For future work, we can explore more robust non-projective SDF sampling methods. Besides, apart from surface reconstruction, our implicit maps with their inherent continuity have the potential to be reused for accurate robot localization [14, 13], which is the focus of our next work.